OpenStudy (anonymous):

Which fractions are greater than -5/16 ? Choose exactly two answers that are correct. A. -3/8 B. - 11/32 C.-1/4 D.-1/8

OpenStudy (anonymous):

A. -3/8 C.-1/4

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

- 5/16 = -0.3125.......... - 3/8 = -.375 and - 11/32= 0.34375

OpenStudy (anonymous):

I'm so sorry its actually A. and B.

OpenStudy (anonymous):

100%

OpenStudy (anonymous):

can you help me with five more?

OpenStudy (anonymous):

@Tyler.Evan2

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

Shoot Them

OpenStudy (anonymous):

k. one second

OpenStudy (anonymous):

.............

OpenStudy (anonymous):

(tick tock tick tock)

OpenStudy (anonymous):

my computer is glitching

OpenStudy (anonymous):

*blah*

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

ok. my questions are really easy, sorry. ill ask you some more later

OpenStudy (anonymous):

oh, ok